OpenStudy (anonymous):

What is the answer to \frac{3}{8}\; \left ( \frac{1}{2} \right )\; +\; \frac{1}{3}? A. 0.4875 B. \frac{5}{16} C. 0.062 D. \frac{25}{48}

OpenStudy (anonymous):

@GabeBae

OpenStudy (gabebae):

This is hard to understand.. i'm sorry

OpenStudy (anonymous):

O. hmmm

OpenStudy (gabebae):

Try taking a screenshot or picture?

OpenStudy (anonymous):

I can do this. \[3/8 What is the answer \to \frac{3}{8}\; \left ( \frac{1}{2} \right )\; +\; \frac{1}{3}? A. 0.4875 B. \frac{5}{16} C. 0.062 D. \frac{25}{48}\]

OpenStudy (anonymous):

Fail

OpenStudy (anonymous):

Sec

OpenStudy (anonymous):

\[ \frac{3}{8}\; \left ( \frac{1}{2} \right )\; +\; \frac{1}{3}?\]

OpenStudy (gabebae):

so you're multiplying and adding fractions?

OpenStudy (anonymous):

I think the question is adding?

OpenStudy (gabebae):

Well, you typed in 3/8 (1/2) I'm assuming you multiply those

OpenStudy (gabebae):

In order to multiply fractions without the same denominator you'll need to find the lowest common denominator

OpenStudy (anonymous):

Nevermind this question. I think I know the answer XD.

OpenStudy (gabebae):

Tell me what the answer is